show please step by step.
595 26825 5301 0001 3240223 558039 121221 eerd MM Va
Answer #1

Solution :

Given that 15,-3,0.52,105,25,-2,0.36,109,18,4,-0.08,115,20,0,1.12,94,23,2,0.5,96,19,-2,-0.16,103,18,3

Mean :-
The mean of a data set is the sum of the terms divided by the total number of terms.

=> Mean μ = sum of terms/number of terms

= 76426/26

= 29.3946

Median :-
The median is the middle number in a sorted list of numbers.

Ordering the data from least to greatest, we get:

-3 -2 -2 -0.16 -0.08 0 0.36 0.5 0.52 1.12 2 3 4 15 18 18 19 20 23 25 94 96 103 105 109 115   

As you can see, we do not have just one middle number but we have a pair of middle numbers, so the median is the average of these two numbers:

=> Median = (4 + 15)/2 = 9.5

Variance :-

The variance of the data set is: σ^2 = 1806.5123

Step 1: Find the mean. [ In this example: =29.3946.(use this calculator for step by step explanation on how to find mean). St

Step 3: Find the sum of numbers in the last column to get. X45162.8076 Step 4: Calculate variance using the above formula. X4

Standard deviation :-

=> The Standard deviation of the data set is σ =

=> Variance σ^2 = 1806.5123

=> Standard deviation = sqrt(Variance)

= sqrt(1806.5123)

= 42.5031